• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fissum

    Log entry time 00:02:03 on November16,2000

    Entry number 52164

    keyword=shift summary


    001115
    16.00 - 24.00 Summary
    KF

    note: this summary has been reconstructed from
    the memories of those on shift for the first two
    hours (see below)

    16.00 - 16.35 Moeller measurement
    16.35 - 16.45 Compton diagnostics
    16.45 - 16.50 FPP chamber 3 trip

    1333.....16.50-17.15.....0.4M.....no endrun event

    17.15 - 18.10 Bad stuff happened. First, ROC14
    died. We went to the tmpwrkDisplay.adl screen and
    toggled crate resets, but the screen we got was a new
    one with no correspondence to the written instructions
    as to how to restart it. Jack returned the screen to its
    old configuration so the instructions once again apply.
    Then we tried a series of kcodas/reboots/runcontrols,
    but the system would not engage. It always hung at
    the et point. Anyway, Bob worked thru it all from home.
    Turned out that whenever the et system started, it core
    dumped. Bob was required to reboot adaqs2 to fix this,
    which clobbered both the Time Account Entry Table
    and the Shift Summary halog.

    1334.....junk

    18.10 MCC took beam for some tweaking

    1335.....18.20-19.22.....1.0M
    1336.....19.22-20.20.....1.0M

    half-wavelength plate removed

    1337.....20.23-21.16.....1.0M
    1338.....21.17-22.07.....1.0M
    1339.....22.08-23.10.....0.9M.....ended run early

    23.10 MCC called. Hardware problems in the south linac.
    Turned out to be a door which blew open and took out the
    interlock.

    1340.....23.16-23.20.....junk

    23.20 Ended 1340 early when the ROCs stopped responding.
    Problem very similar to earlier this evening. Reboots,
    kcodas, etc. ROC2 and ROC3 in particular are very, very
    stubborn.

    1341.....23.33-